can someone PLEASE help me solve this system:
{ x+ y-4z=6
{2x-5y+3z=2
{ x-6y+7z=-4
thank you so much!

Answer by stanbon(75887)

x+ y -4z=6
2x-5y+3z=2
x-6y+7z=-4
---------------
Using the TI-83 Matrix function I find
the determinant of the coefficient matrix
is zero. So there is no solution for this
system of equations.
-------------------
You can see why this happens if you add
equation #1 to equation #3. You get
2x-5y+3 = 2. Therefore the system of
equation is dependent; the three planes
are really just one plane.
--------------------
Cheers,
Stan H.